# Probing Picohertz Gravitational Waves with Pulsars

### v1

@@ Appendix A Derivation of analytic model
where $\Psi$,$i$ defines the orientation of the GW source, and $\phi=\phi_{0}+\omega_{0}t$, with $\omega_{0}$ the orbital frequency of the binary gravitational wave source.
@@
$v_{GW}=\frac{\hat{d_{p}^{i}}\hat{d_{p}^{j}}}{2(1+\hat{\Omega}\cdot\hat{d_{p}})}[e_{ij}^{+}(\hat{\Omega})(h_{0}\cos(2\Psi)(1+\cos^{2}(i))\sin(2\phi_{0}+2\omega_{0}t)+2h_{0}\sin(2\Psi)\cos(i)\cos(2\phi_{0}+2\omega_{0}t)-h_{0}\cos(2\Psi)(1+\cos^{2}(i))\sin(2\phi_{0}+2\omega_{0}(t-\hat{\Omega}\cdot\vec{d_{p}}))-2h_{0}\sin(2\Psi)\cos(i)\cos(2\phi_{0}+2\omega_{0}(t-\hat{\Omega}\cdot\vec{d_{p}})))+e_{ij}^{\times}(\hat{\Omega})(h_{0}\sin(2\Psi)(1+\cos^{2}(i))\sin(2\phi_{0}+2\omega_{0}t)-2h_{0}\cos(2\Psi)\cos(i)\cos(2\phi_{0}+2\omega_{0}t)-h_{0}\sin(2\Psi)(1+\cos^{2}(i))\sin(2\phi_{0}+2\omega_{0}(t-\hat{\Omega}\cdot\vec{d_{p}}))+2h_{0}\cos(2\Psi)\cos(i)\cos(2\phi_{0}+2\omega_{0}(t-\hat{\Omega}\cdot\vec{d_{p}})))]$
The direction of the source can be parametrized by the galactic longitude $l$ and latitude $b$. In the Cartesian coordinates,
$\begin{cases}\hat{l}=-\sin(l)\hat{i}+\cos(l)\hat{j}\\ \hat{b}=\cos(b)\cos(l)\hat{i}+\cos(b)\sin(l)\hat{j}-\sin(b)\hat{k}\end{cases}$ (11)
@@
$\begin{cases}e_{+}=\hat{l}\hat{l}-\hat{b}\hat{b}\\ e_{\times}=\hat{l}\hat{b}+\hat{b}\hat{l}\end{cases}$ (12)
In the analysis, we need to use the acceleration and jerk $a_{GW}$, $j_{GW}$, which we derive here. The jerk is easy to derive as it only pulls out a factor of $(2\omega_{0})^{2}$ in front of the original expression for $v_{\text{GW}}$. The acceleration requires a little more work where one takes the derivatives of each term carefully. Because we are working in the pHz frequency regime with typical periods much longer than the observation time, we can safely assume that $\omega_{0}t<<1$, and thus expand $\cos(2\omega_{0}t)$ as 1, and $\sin(2\omega_{0}t)$ as $2\omega_{0}t$. Now we have
$a_{GW}=\frac{\hat{d_{p}^{i}}\hat{d_{p}^{j}}}{1+\hat{\Omega}\cdot\hat{d_{p}}}\omega_{0}h_{0}[e_{ij}^{+}(\hat{\Omega})[\cos(2\Psi)(1+\cos^{2}(i))(\cos(2\phi_{0})-\cos(2\phi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))+2\sin(2\Psi)\cos(i)(-\sin(2\phi_{0})+\sin(2\phi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))]+e_{ij}^{\times}(\hat{\Omega})[\sin(2\Psi)(1+\cos^{2}(i))(\cos(2\phi_{0})-\cos(2\phi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))-2\cos(2\Psi)\cos(i)(-\sin(2\phi_{0})+\sin(2\phi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))]]$ (13)
$j_{GW}=-2\frac{\hat{d_{p}^{i}}\hat{d_{p}^{j}}}{1+\hat{\Omega}\cdot\hat{d_{p}}}\omega_{0}^{2}h_{0}[e_{ij}^{+}(\hat{\Omega})[\cos(2\Psi)(1+\cos^{2}(i))(\sin(2\phi_{0})-\sin(2\phi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))+2\sin(2\Psi)\cos(i)(\cos(2\phi_{0})-\cos(2\phi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))]+e_{ij}^{\times}(\hat{\Omega})[\sin(2\Psi)(1+\cos^{2}(i))(\sin(2\phi_{0})-\sin(2\phi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))-2\cos(2\Psi)\cos(i)(\cos(2\phi_{0})-\cos(2\phi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))]]$ (14)
The unit vectors $\hat{d_{p}}$ and $\hat{\Omega}$ can be parametrized by the equatorial ascension($\alpha_{p},l$) and declination ($\delta_{p},b$) of the pulsars and source in the Cartesian coordinates, which read
$\begin{cases}\hat{d_{p}}=\sin(\delta_{p})\cos(\alpha_{p})\hat{i}+\sin(\delta_{p})\sin(\alpha_{p})\hat{j}+\cos(\delta_{p})\hat{k}\\ \hat{\Omega}=-\sin(b)\cos(l)\hat{i}+\sin(b)\sin(l)\hat{j}+\cos(b)\hat{k}\end{cases}$ (15)
@@ Appendix C Pulsar data
J0030+0451 113.141 -57.611 0.324 4(4)-4(4) 0.54 [37]
J0034-0534 111.492 -68.069 1.348 $0(20)$ 0 [37]
J0218+4232 139.508 -17.527 3.150 2(5)-2(5) 0.14 [37]
J0437-4715 253.394 -41.963 0.157 1(1)-1(1) 0 [38]
J0610-2100 227.747 -18.184 3.260 $0(50)$ 0 [37]
J0613-0200 210.413 -9.305 0.780 $0.6(6)$ 0.13 [37]
J0621+1002 200.570 -2.013 0.425 70(30)-70(30) 1.28 [37]
J0711-6830 279.531 -23.280 0.106 $1(1)$ 0 [38]
J0751+1807 202.730 21.086 1.110 $0(2)$ 0 [37]
J0900-3144 256.162 9.486 0.890 10(20)-10(20) 0 [37]
J1012+5307 160.347 50.858 0.700 $0.4(7)$ 0.01 [37]
J1022+1001 231.795 51.101 0.645 2(1)-2(1) 0.01 [37]
J1045-4509 280.851 12.254 0.340 2(7)-2(7) 0.05 [38]
J1455-3330 330.722 22.562 0.684 $6(20)$ 0.17 [37]
J1600-3053 344.090 16.451 1.887 $4(5)$ 0.06 [38]
J1603-7202 316.630 -14.496 0.530 $1(4)$ 0.02 [38]
J1640+2224 41.051 38.271 1.515 0.9(9)-0.9(9) 0 [37]
J1643-1224 5.669 21.218 0.740 2(2)-2(2) 0.01 [37]
J1713+0747 28.751 25.223 1.311 0.5(5)-0.5(5) 0.38 [37]
J1721-2457 8.267 6.751 1.393 30(70)-30(70) 0.01 [37]
J1730-2304 3.137 6.023 0.620 $0(2)$ 0 [38]
J1732-5049 340.029 -9.454 1.873 $20(20)$ 0.03 [38]
J1738+0333 27.721 17.742 1.471 30(90)-30(90) 0 [37]
J1744-1134 14.794 9.180 0.395 $0.8(8)$ 0.02 [37]
J1751-2857 0.646 -1.124 1.087 10(50)-10(50) 0 [37]
J1801-1417 14.546 4.162 1.105 30(100)-30(100) 0.02 [37]
J1802-2124 8.382 0.611 0.760 $10(60)$ 0.01 [37]
J1804-2717 3.508 -2.736 0.805 40(40)-40(40) 0 [37]
J1843-1113 22.055 -3.397 1.260 7(20)-7(20) 0.05 [37]
J1853+1303 44.875 5.367 2.083 30(20)-30(20) 0 [37]
B1855+09 42.290 3.060 1.200 $1(2)$ 0.03 [37]
J1909-3744 359.731 -19.596 1.140 $0.6(9)$ 0.02 [37]
J1910+1256 46.564 -1.795 1.496 $30(20)$ 0 [37]
J1911+1347 25.137 -9.579 1.069 $14(8)$ 0 [37]
J1911-1114 47.518 1.809 1.365 $20(50)$ 0 [37]
J1918-0642 30.027 -9.123 1.111 $0(8)$ 2.46 [37]
B1953+29 65.839 0.443 6.304 20(50)-20(50) 0 [37]
J2010-1323 29.446 -23.540 2.439 $20(20)$ 0 [37]
J2019+2425 64.746 -6.624 1.163 500(900)-500(900) 0 [37]
J2033+1734 60.857 -13.154 1.740 $40(100)$ 0 [37]
J2124-3358 10.925 -45.438 0.410 $0(3)$ 0.02 [38]
J2129-5721 338.005 -43.570 3.200 1(2)-1(2) 0 [38]
J2145-0750 47.777 -42.084 0.714 2(1)-2(1) 0.28 [37]
J2229+2643 87.693 -26.284 1.800 20(20)-20(20) 0 [37]
J2317+1439 91.361 -42.360 1.667 1(3)-1(3) 0 [37]
J2322+2057 96.515 -37.310 1.011 $30(70)$ 0 [37]
Table 2: Pulsar timing parameters including galactic coordinates, distance ($d$), fractional spin-down rate $\ddot{P}_{\text{obs}}/P$, timing noise strength $\sigma_{\text{RN}}$, and references.
Pulsar $l$ (deg) $b$ (deg) d (kpc) $\Delta\dot{P}/P_{b}(10^{-18}s^{-1})$
B1259-63 304.184 -0.992 2.60(35) 130.7(654)
B1534+12 19.848 48.341 0.940(65) 0.203(109)
B1913+16 49.968 2.122 4.10(135) -1.256(83)
J0348+0432 183.337 -36.774 1.6(9) -1.82(608)
J0437-4715 253.394 -41.963 0.156(11) 0.0508(555)
J0613-0200 210.413 -9.305 1.11(5) -0.10(18)
J0737-3039A/B 245.236 -4.505 0.735(60) -0.475(3755)
J0740+6620 149.730 29.599 1.15(16) -0.04(64)
J0751+1807 202.730 21.086 1.40(35) -0.326(967)
J1012+5307 160.347 50.858 0.845(14) 0.164(229)
J1017-7156 291.558 -12.553 1.4(18) 0.518(571)
J1022+1001 231.795 51.101 0.850(55) -0.081(110)
J1125-6014 292.504 0.894 1.5(8) -0.017(586)
J1455-3330 330.722 22.562 0.760(55) 0.608(337)
J1537+1155 19.850 48.340 1.16(24) -0.09(38)
J1600-3053 344.090 16.451 1.39(4) 0.181(82)
J1603-7202 316.630 -14.496 0.53(28) 0.293(84)
J1614-2230 352.636 20.192 0.690(45) 0.315(218)
J1640+2224 41.051 38.271 1.08(24) 0.439(153)
J1713+0747 28.751 25.223 1.31(4) -0.026(22)
J1738+0333 27.721 17.742 1.74(14) 0.028(344)
J1741+1351 37.885 21.641 1.8(4) 22.6(63)
J1909-3744 359.731 -19.596 1.14(1) 0.049(34)
J1933-6211 334.431 -28.632 1.2(4) -0.336(143)
J2043+1711 61.919 -15.313 1.6(2) 107.97(6481)
J2129-5721 338.005 -43.570 0.6(4) 2.51(24)
J2145-0750 47.777 -42.084 0.62(2) 0.11(5)
J2222-0137 62.018 -46.075 0.268100(105) 0.000098(45995)
J2234+0611 72.991 -43.006 0.95(4) -0.36(90)
J2339-0533 81.349 -62.476 1.44(30) 8.01(525)
Table 3: Complete table of pulsar data including Galactic coordinates, distance, and fractional change in orbital period. [34] The contribution from galactic acceleration is calculated using galpy [35].

### v2

@@ Appendix A Derivation of analytic model
where $\phi$, $i$ defines the orientation of the GW source, and $\Psi=\Psi_{0}+\omega_{0}t$, with $\omega_{0}$ the orbital frequency of the binary gravitational wave source.
@@
$v_{GW}=\frac{\hat{d_{p}^{i}}\hat{d_{p}^{j}}}{2(1+\hat{\Omega}\cdot\hat{d_{p}})}[e_{ij}^{+}(\hat{\Omega})(h_{0}\cos(2\phi)(1+\cos^{2}(i))\sin(2\Psi_{0}+2\omega_{0}t)+2h_{0}\sin(2\phi)\cos(i)\cos(2\Psi_{0}+2\omega_{0}t)-h_{0}\cos(2\phi)(1+\cos^{2}(i))\sin(2\Psi_{0}+2\omega_{0}(t-\hat{\Omega}\cdot\vec{d_{p}}))-2h_{0}\sin(2\phi)\cos(i)\cos(2\Psi_{0}+2\omega_{0}(t-\hat{\Omega}\cdot\vec{d_{p}})))+e_{ij}^{\times}(\hat{\Omega})(h_{0}\sin(2\phi)(1+\cos^{2}(i))\sin(2\Psi_{0}+2\omega_{0}t)-2h_{0}\cos(2\phi)\cos(i)\cos(2\Psi_{0}+2\omega_{0}t)-h_{0}\sin(2\phi)(1+\cos^{2}(i))\sin(2\Psi_{0}+2\omega_{0}(t-\hat{\Omega}\cdot\vec{d_{p}}))+2h_{0}\cos(2\phi)\cos(i)\cos(2\Psi_{0}+2\omega_{0}(t-\hat{\Omega}\cdot\vec{d_{p}})))]$
The direction of the source can be parametrized by the right ascension $\alpha$ and declination $\delta$. In the Cartesian coordinates,
$\begin{cases}\hat{\alpha}=-\sin(\alpha)\hat{i}+\cos(\alpha)\hat{j}\\ \hat{\delta}=\cos(\delta)\cos(\alpha)\hat{i}+\cos(\delta)\sin(\alpha)\hat{j}-\sin(\delta)\hat{k}\end{cases}$ (11)
@@
$\begin{cases}e_{+}=\hat{\alpha}\hat{\alpha}-\hat{\delta}\hat{\delta}\\ e_{\times}=\hat{\alpha}\hat{\delta}+\hat{\delta}\hat{\alpha}\end{cases}$ (12)
In the analysis, we need to use the acceleration and jerk $a_{GW}$, $j_{GW}$, which we derive here. Since we are working in the pHz frequency range with typical periods much longer than the observation time, we can safely assume that $\omega_{0}t<<1$ in our derivations, and thus expand $\cos(2\omega_{0}t)$ as 1, and $\sin(2\omega_{0}t)$ as $2\omega_{0}t$. Now we have
$a_{GW}=\frac{\hat{d_{p}^{i}}\hat{d_{p}^{j}}}{1+\hat{\Omega}\cdot\hat{d_{p}}}\omega_{0}h_{0}[e_{ij}^{+}(\hat{\Omega})[\cos(2\phi)(1+\cos^{2}(i))(\cos(2\Psi_{0})-\cos(2\Psi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))+2\sin(2\phi)\cos(i)(-\sin(2\Psi_{0})+\sin(2\Psi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))]+e_{ij}^{\times}(\hat{\Omega})[\sin(2\phi)(1+\cos^{2}(i))(\cos(2\Psi_{0})-\cos(2\Psi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))-2\cos(2\phi)\cos(i)(-\sin(2\Psi_{0})+\sin(2\Psi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))]]$ (13)
$j_{GW}=-2\frac{\hat{d_{p}^{i}}\hat{d_{p}^{j}}}{1+\hat{\Omega}\cdot\hat{d_{p}}}\omega_{0}^{2}h_{0}[e_{ij}^{+}(\hat{\Omega})[\cos(2\phi)(1+\cos^{2}(i))(\sin(2\Psi_{0})-\sin(2\Psi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))+2\sin(2\phi)\cos(i)(\cos(2\Psi_{0})-\cos(2\Psi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))]+e_{ij}^{\times}(\hat{\Omega})[\sin(2\phi)(1+\cos^{2}(i))(\sin(2\Psi_{0})-\sin(2\Psi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))-2\cos(2\phi)\cos(i)(\cos(2\Psi_{0})-\cos(2\Psi_{0}-2\omega_{0}(1+\hat{\Omega}\cdot\hat{d_{p}})d_{p}))]]$ (14)
The unit vectors $\hat{d_{p}}$ and $\hat{\Omega}$ can be parametrized by the equatorial ascension($\alpha_{p},\alpha$) and declination ($\delta_{p},\delta$) of the pulsars and source in the Cartesian coordinates, which read
$\begin{cases}\hat{d_{p}}=\sin(\delta_{p})\cos(\alpha_{p})\hat{i}+\sin(\delta_{p})\sin(\alpha_{p})\hat{j}+\cos(\delta_{p})\hat{k}\\ \hat{\Omega}=-\sin(\delta)\cos(\alpha)\hat{i}+\sin(\delta)\sin(\alpha)\hat{j}+\cos(\delta)\hat{k}\end{cases}$ (15)
@@ Appendix C Pulsar data
Pulsar $l$ (deg) $b$ (deg) d (kpc) $\Delta\dot{P}_{b}/P_{b}(10^{-18}s^{-1})$
B1259-63 304.184 -0.992 2.6(4) 130(70)
B1534+12 19.848 48.341 0.94(7) 0.2(1)
B1913+16 49.968 2.122 4(1) -1.26(8)
J0348+0432 183.337 -36.774 1.6(9) -2(6)
J0437-4715 253.394 -41.963 0.16(1) 0.05(6)
J0613-0200 210.413 -9.305 1.11(5) -0.1(2)
J0737-3039A/B 245.236 -4.505 0.74(6) -0.5(40)
J0740+6620 149.730 29.599 1.2(2) -0.04(60)
J0751+1807 202.730 21.086 1.4(4) -0.3(10)
J1012+5307 160.347 50.858 0.85(1) 0.2(2)
J1017-7156 291.558 -12.553 1(2) 0.5(6)
J1022+1001 231.795 51.101 0.85(6) -0.08(11)
J1125-6014 292.504 0.894 1.5(8) -0.02(60)
J1455-3330 330.722 22.562 0.76(6) 0.6(3)
J1537+1155 19.850 48.340 1.2(2) -0.09(40)
J1600-3053 344.090 16.451 1.39(4) 0.18(8)
J1603-7202 316.630 -14.496 0.5(3) 0.29(8)
J1614-2230 352.636 20.192 0.69(5) 0.3(2)
J1640+2224 41.051 38.271 1.0(2) 0.4(2)
J1713+0747 28.751 25.223 1.31(4) -0.03(2)
J1738+0333 27.721 17.742 1.7(1) 0.03(30)
J1741+1351 37.885 21.641 1.8(4) 23(6)
J1909-3744 359.731 -19.596 1.14(1) 0.05(3)
J1933-6211 334.431 -28.632 1.2(4) -0.3(1)
J2043+1711 61.919 -15.313 1.6(2) 107(60)
J2129-5721 338.005 -43.570 0.6(4) 2.5(2)
J2145-0750 47.777 -42.084 0.62(2) 0.11(5)
J2222-0137 62.018 -46.075 0.2681(1) 0.00(5)
J2234+0611 72.991 -43.006 0.95(4) -0.4(9)
J2339-0533 81.349 -62.476 1.4(3) 8(5)
Table 2: Complete table of pulsar data including Galactic coordinates, distance, and fractional acceleration in orbital period $\Delta\dot{P}_{b}/P_{b}$. This quantity includes all contributions that should sum up to $a_{\text{GW}}$ as in Eq. 3. Uncertainties are reported as 1 standard deviation. The contribution from Galactic acceleration is calculated using galpy [Bovy2015], and all distances are measured using parallax. The distance and timing parameter data can be found in [Moran:2023, DeRoccoDror2024]. The pulsars satisfying the distance measurement criterion discussed in Sec. IV are in bold face.
@@
J0030+0451 113.141 -57.611 0.3(1)[Arzoumanian_2018] 4(4)-4(4) 0.54 [Desvignes2016]
J0034-0534 111.492 -68.069 1.3(5)[yao2017new] $0(20)$ 0 [Desvignes2016]
J0218+4232 139.508 -17.527 3(1)[Verbiest_2014] 2(5)-2(5) 0.14 [Desvignes2016]
J0437-4715 253.394 -41.963 0.16(1)[Moran:2023] 1(1)-1(1) 0 [Reardon2016]
J0610-2100 227.747 -18.184 3(1)[yao2017new] $0(50)$ 0 [Desvignes2016]
J0613-0200 210.413 -9.305 1.11(5)[Moran:2023] $0.6(6)$ 0.13 [Desvignes2016]
J0621+1002 200.570 -2.013 0.4(2)[yao2017new] 70(30)-70(30) 1.28 [Desvignes2016]
J0711-6830 279.531 -23.280 0.11(4)[yao2017new] $1(1)$ 0 [Reardon2016]
J0751+1807 202.730 21.086 1.4(4)[Moran:2023] $0(2)$ 0 [Desvignes2016]
J0900-3144 256.162 9.486 0.8(4)[Desvignes2016] 10(20)-10(20) 0 [Desvignes2016]
J1012+5307 160.347 50.858 0.85(1)[Moran:2023] $0.4(7)$ 0.01 [Desvignes2016]
J1022+1001 231.795 51.101 0.85(6)[Moran:2023] 2(1)-2(1) 0.01 [Desvignes2016]
J1045-4509 280.851 12.254 0.3(2)[Bobakov_2019] 2(7)-2(7) 0.05 [Reardon2016]
J1455-3330 330.722 22.562 0.76(6)[Moran:2023] $6(20)$ 0.17 [Desvignes2016]
J1600-3053 344.090 16.451 1.39(4)[Moran:2023] $4(5)$ 0.06 [Reardon2016]
J1603-7202 316.630 -14.496 0.5(3)[Moran:2023] $1(4)$ 0.02 [Reardon2016]
J1640+2224 41.051 38.271 1.0(2)[Moran:2023] 0.9(9)-0.9(9) 0 [Desvignes2016]
J1643-1224 5.669 21.218 0.8(2)[Desvignes2016] 2(2)-2(2) 0.01 [Desvignes2016]
J1713+0747 28.751 25.223 1.31(4)[Moran:2023] 0.5(5)-0.5(5) 0.38 [Desvignes2016]
J1721-2457 8.267 6.751 1.4(6)[yao2017new] 30(70)-30(70) 0.01 [Desvignes2016]
J1730-2304 3.137 6.023 0.44(5)[Shamohammadi:2024uvg] $0(2)$ 0 [Reardon2016]
J1732-5049 340.029 -9.454 1.9(7)[yao2017new] $20(20)$ 0.03 [Reardon2016]
J1738+0333 27.721 17.742 1.7(1)[yao2017new] 30(90)-30(90) 0 [Desvignes2016]
J1744-1134 14.794 9.180 0.38(1)[Shamohammadi:2024uvg] $0.8(8)$ 0.02 [Desvignes2016]
J1751-2857 0.646 -1.124 1.1(4)[yao2017new] 10(50)-10(50) 0 [Desvignes2016]
J1801-1417 14.546 4.162 2(1)[Shamohammadi:2024uvg] 30(100)-30(100) 0.02 [Desvignes2016]
J1802-2124 8.382 0.611 0.6(4)[Desvignes2016] $10(60)$ 0.01 [Desvignes2016]
J1804-2717 3.508 -2.736 0.8(3)[yao2017new] 40(40)-40(40) 0 [Desvignes2016]
J1843-1113 22.055 -3.397 5(3)[gaia_2023] 7(20)-7(20) 0.05 [Desvignes2016]
J1853+1303 44.875 5.367 0.9(4)[Desvignes2016] 30(20)-30(20) 0 [Desvignes2016]
B1855+09 42.290 3.060 0.9(2)[Verbiest_2012] $1(2)$ 0.03 [Desvignes2016]
J1909-3744 359.731 -19.596 1.14(1)[Moran:2023] $0.6(9)$ 0.02 [Desvignes2016]
J1910+1256 46.564 -1.795 2.5(5)[K_t_kc__2022] $30(20)$ 0 [Desvignes2016]
J1911+1347 25.137 -9.579 2.0(3)[K_t_kc__2022] $14(8)$ 0 [Desvignes2016]
J1911-1114 47.518 1.809 1.1(4)[yao2017new] $20(50)$ 0 [Desvignes2016]
J1918-0642 30.027 -9.123 1.2(3)[Shamohammadi:2024uvg] $0(8)$ 2.46 [Desvignes2016]
B1953+29 65.839 0.443 6(3)[yao2017new] 20(50)-20(50) 0 [Desvignes2016]
J2010-1323 29.446 -23.540 3(2)[Shamohammadi:2024uvg] $20(20)$ 0 [Desvignes2016]
J2019+2425 64.746 -6.624 1.2(5)[yao2017new] 500(900)-500(900) 0 [Desvignes2016]
J2033+1734 60.857 -13.154 1.7(7)[yao2017new] $40(100)$ 0 [Desvignes2016]
J2124-3358 10.925 -45.438 0.48(4)[Shamohammadi:2024uvg] $0(3)$ 0.02 [Reardon2016]
J2129-5721 338.005 -43.570 0.6(4)[Moran:2023] 1(2)-1(2) 0 [Reardon2016]
J2145-0750 47.777 -42.084 0.62(2)[Moran:2023] 2(1)-2(1) 0.28 [Desvignes2016]
J2229+2643 87.693 -26.284 1.8(7)[yao2017new] 20(20)-20(20) 0 [Desvignes2016]
J2317+1439 91.361 -42.360 2.0(4)[Arzoumanian_2018] 1(3)-1(3) 0 [Desvignes2016]
J2322+2057 96.515 -37.310 0.8(2)[Shamohammadi:2024uvg] $30(70)$ 0 [Desvignes2016]
Table 3: Pulsar timing parameters including Galactic coordinates, distance, fractional spin-down acceleration $\ddot{P}_{\text{obs}}/P$ as in Eq. 4, variance induced by red noise $\sigma_{\text{RN}}$, and references. The uncertainties are reported as 1 standard deviation. For pulsars without a parallax measurement, the electron density model YMW 17 [yao2017new] is used to estimate the distance. A relative uncertainty of $40\%$ is assumed in such cases. References for both the pulsar distances and the timing parameter data are included.